FEL Upgrade Project June 2002 Monthly Report



This month the installation crew held the Upgrade’s first "grouting" party. With the assistance of many of the FEL team members who dressed for their part, all of the stands which had previously been surveyed and installed for FEL Upgrade hardware were grouted to the floor.

We warmed up the injector cryounit this month so that we could replace the rf windows for the higher powers required for the 10 kW upgrade, replace a leaky valve, and check the alignment prior to installing new gun and injector hardware.

The last major procurement for the IR Upgrade, the large vacuum chambers for the recirculation arcs went out to bid this month. The measurements on the test stand for the last of the quadrupole magnets were completed. Thus all the quads required for both the IR and UV upgrades are now available for installation.

We had a very productive final week of the month: (1) the RF system for the second cryomodule (Zone 4) was completely checked out except for the final SRF test when the 2nd cryomodule is installed. All 8 klystrons are operational; (2) the new High Voltage tank for the Upgrade gun arrived on site and was installed in the FEL vault; (3) all the quadrupole girder assemblies between the first arc and the optical chicane are installed;(4) we received favorable bids on our last major procurement (the large arc vacuum chambers) and placed the order early in July.


As noted in the financial report section below, both the IR and UV Upgrade projects are on track; however, because JLab is still due $1.5M of AF funding and $1.2M of ONR/JTO funding for FY02, there are budget and cash flow problems building as we approach the end of the fiscal year. We have asked for the assistance of several folks in our supply chain to help solve these funding release problems. We thank Gil Graff, John Eric and Wayne Skinner for their assistance in these matters.

We prepared dry run presentations for our upcoming annual DOE review of Jefferson Lab’s Science and Technology scheduled for July 15-16.

The review process continues. The draft report from the May 1-3 semiannual review of the FEL Upgrade project was reviewed by the project team and sent back to the ONR contract monitor for distribution. We thank the panel for their constructive and fair review of our progress.

We attended two separate meetings of JTO Technical Working Groups: the lethality working group met on the 18th at the Dahlgren NSWC and the FEL working group met on the 21st at the University of Maryland.

On Monday, June 10, F. Dylla met with the management of the Wellman Laboratory of Photomedicine at Massachusetts General Hospital in Boston to continue our discussions for collaborative efforts on FEL applications to photobiology and photomedicine.


On Tuesday, June 18th, G. Neil attended a meeting of the various groups that are participating in laser lethality studies funded by the JTO. The meeting was held at the Navy’s Dahlgren Labs. Our plan for offering the FEL User facility as a nationwide resource for continuing JTO (and other DOD funded) lethality studies is gaining support. We plan on continuing our partnership with NRL and MIT Lincoln Labs with these studies.

On Tuesday, June 18th, F. Dylla and G. Williams attended a meeting at UVA to continue our discussions on use of the FEL for materials science and biology/biomedical studies and planning for a collaboration in the use of advanced light sources.

On Thursday, June 20th, we had visits from Major Steve Redifer, who will become the Marine Corp. representative on directed energy development. Major Redifer is currently finishing graduate studies at the NPS with our colleague, Prof. Bill Colson. Also on Thursday we had a meeting with Alan Todd and Tony Favale from AES, Inc. to discuss our progress on the CRADA for the development of the 100 mA injector.

On Thursday June 27, F. Dylla attended the quarterly meeting of the Virginia Research and Technology Advisory Commission which is chaired by RADM (ret) Marc Pelaez and Prof. Anita Jones (UVA). The VRTAC has always been very supportive of the FEL project.

We met this month with Brian Hankla of the JTO Lethality TAWG to discuss continuation of our FEL laser effects testing program on materials. Our proposal to construct a dedicated suite of materials diagnostics and to continue with beam tests has a high priority with the Navy. We expect about $600k to be provided for this effort next year.

Project Cost Performance:

The Phase 1 FEL Upgrade project budget for the period June 1, 2000 to September 30, 2001 is $9,029k. The first phase of the project was fully obligated as of the end of September 2001. The total Phase 1 project cost to date-actual costs plus commitments is $9,029k as planned for Phase 1 completion.

The Phase 2 FEL Upgrade project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002. Actual funds transferred from ONR to DOE for this effort in FY02 were $4,442k. Work for Phase 2 proceeded to plan during this reporting period. Actual charges of $361k were accrued for June for total accrued costs of $3,637k to date for Phase 2. Work scheduled for June was $710k for a total of $3,389k for Phase 2 to date; work performed for June was $479k for a total of $3,274k for Phase 2 to date, which results in 76% scheduled vs. 74% performed for the Phase 2 project. At present the cost variance for Phase 2 is -$363k and the schedule variance is -$115k. We expect to meet the planned cost at completion of Phase 2 effort. The program is presently on track to begin commissioning efforts in September 2002.

The UV FEL project was approved for $2,836k for a performance period from September 2001 to September 30, 2002. Actual charges of $467k were accrued through June for a total accrued costs of $1,805k since project start. Work scheduled for June was $286k for a total of $1,701k for the UV project to date, work performed for June was $220k for a total of $1,699k for the UV project to date, which results in 60% scheduled vs. 60% performed for the UV project. At present the cost variance for the UV project is -$107k and the schedule variance is -$2k. The program is on track to meet the statement of work requirements on schedule and within the available budget.

WBS 3 (Beam Physics):

Initial (organizational) planning and detailing of the commissioning plan started. Design of a very compact, very large acceptance energy recovery system continued.

WBS 4 (Injector):

The high voltage power supply in the W&M implantation has been repaired after the high voltage connector failed. A test run with a sample electrode inside the plasma proved that the system is working properly now. The shielding door for the ball cathode was successfully implanted this month. The support tube which suffered arc damage in the last run has had the implant layer etched away and is being polished.

A HEPA filter unit and new curtains have been installed around the W&M implant system creating a much cleaner area. The cables on the system were replaced with shielded cables to reduce noise. The ball cathode requires additional polishing and is being worked.

The tapered shielded bellows drawings are in sign off, and the bellows for it have been received. The shielded bellows/BPM unit has had the BPM tested and the design is now finalized, ordered bellows for the unit. The SF6 tank modification drawings are in review.

The gun stand and chamber were installed in the injector pit and is being aligned

The final drawings for the light box support structure were signed off this month. The drawing for the light box transport was ready for final check. The machine shop is making good progress on the light box can. The only design work remaining on the light box is the alignment drawing.

Gun HVPS - The Gun HVPS Conditioning Resistor was fired this month. After looking at the results and then sending photo's to the vendor of the resistive material, it was determined the resistor was at the firing temperature for too long a time. It is not possible to reduce the firing time, so a lower temperature was suggested. The resistive material formed small, unconnected puddles on the silver contacts. The resistor section looked like an open circuit above 1 kM. This resistor will be used to determine a better firing profile. It was physically broken during handling, but it was already unusable due to the over-firing. The next firing will be done at a lower temperature.

The tank for the Gun's HVPS was delivered this month and all but the door was installed. The Gun Tank must be installed before the final mounting alignment can be completed and tank's door is installed. Due to dust and dirt in the vault getting on the HV components, the HVPS will be installed after the door is in place and can be closed. The draft SOP was sent out for comments and any corrections.

WBS 6 (RF):

Quarter Klystrons - The klystrons and their solenoids for the Quarter were installed on their roll-around carts this week. The plumbing and wiring will be finished as time permits.

Quarter HVPS - The PLC code to control the klystron heater and the SCR controller was successfully tested on unit 3. The printed circuit board to monitor the heater voltage was released this week for fabrication. The draft SOP is in progress and will be sent out for

comments and corrections next week.


Zone 4 - The zone was restored to operating condition. RF was run through all 8 klystrons. The RF system for unit 4-8 was calibrated and data curves recorded. All systems look good and are ready for the cryomodule. The MPS & PSS equipment needs to be installed for this zone. Prior to SRF checkout of the second cryomodule that is being installed in July, the MPS and PSS systems need to be activated and certified again. The second FEL cryomodule (returning from SL21in CEBAF) is expected to be installed on July 9th with a cooldown on July 24th.


WBS 8 (Instrumentation):

Progress is made on the interface for the beam viewer chassis, assembly of Beam Loss Monitor heads, and fabrication of 8 channel stepper motor drives.

Production issues were worked through with assembly of quad girders. The short stripline BPM for the injector was tested after welding in of the feedthroughs. There was a problem with one of the four, it is being repaired.

The last magnet trim rack feeds have been cabled into the junction box downstairs in the vault. A 3-phase transformer and breaker panel for the magnet power supplies has been installed in the west end of the gallery. This will simplify the AC power connections to the new magnet supplies and enhance safety by providing a single point to lock and tag the new power supplies.

The final cabling is being done for the girders installed in the tunnel. All the light rings for the

beam viewers have been stuffed and tested. Camera assemblies for MQX3F04, MQX3F05, MQX3F06 have been installed.



Work continued on the fabrication for the Beam Loss Monitors. The parts for the Photo Multiplier Tube (PMT) are all in and are being assembled by one of our students. The new Canbus Interface board for the beam viewer control chassis upgrade has been completed and is at the board house with a delivery date of 10 days.

Six of the BLM rear board prototypes (F0150) have been sent out for fabrication to be received on a 5-day turn. Front panels for the boards have been ordered as well and all the required parts are on order. Final adjustments to the BLM front card (F0151) schematic have been made and the package is now back with EECAD. Prototypes are planed to go out for fabrication in early July.

Additional improvements have been made to the ion implantation system. A remote control box is being discussed such that the metal rack door could be kept closed during processing to help shield electrical noise from the Pulse Forming Network (PFN).

Progress has been made on the drive laser pulse control system. The design has progressed to where the ConOptics Model 305 (Synchronous Countdown Chassis which we have used for micropulse control) eliminated by placing the core logic board on to a VME carrier for better integration into the control system. The 6U VME board is being designed to support & power the board. The motivation is to package the entire system in a single crate and eliminate the many cables and interfaces to the various laser control chassis. Design work also continues on a board that monitors and interlocks the external RF to the drive laser. The external oscillator for the mode locker will also be mounted to a VME board.

A meeting was held to discuss the requirements for the 100 amp sextapole/octapole power supplies. The issues were whether we could take credit for damping of the ripple voltage since they are solid core magnets. The plan is to move forward with a low cost power supply and purchase an external current transductor. The piece price of the power supplies is ~2500 + 1000 for the transductor vs ~$15k for a 10PPM supply. If the resulting magnetic field ripple is too high (as measured on test stand) the supplies will be fitted with the current transductor and current loop regulator with sufficient gain at frequencies 60 Hz etc. The controls for this approach will rely on same devices as for the super trim supplies. This is the SNS/BNL fiber isolated controller (EPICS software is already complete!).

The front panel pcb schematic for the OMS chassis (8 channel stepper motor driver) has been completed and submitted to EECAD. The fast raster control board design is complete and submitted to EECAD, this is used to create the recirculating dump raster pattern. Power cables for the first 32 channels of trim magnets for the Injector (MFF0F01 thru MQX1F03) have been pulled into place. The rework of the injector cables continues.

Nate (our DOE student) has completed the gain and characterizing of the BLM's (good job!). All assemblies have been tested, packed and stored. Additionally he investigated response to various other wavelength LEDs.

Work continues on the fast raster for the recirculating beam dump. The system consists of a control chassis upstairs in the control room and a tunnel driver chassis near the magnet. Three (horizontal & vertical + spare) of the tunnel boards have been stuffed and are currently being tested. The upstairs Fast Raster board check prints have been turned in to EECAD for final sign off. The new PCBs should be ready for testing in a couple weeks.

The Canbus beam viewer interface boards have been received from the board house and the CAD drawings turned over to the EECAD. These will stuffed (surface mount) and tested then sent to Al Grippo for software debugging.

The students are working on all of the girder connections including air and water. The beam viewer assemblies, cameras and lamp rings on the girders have been installed and tested up to the viewer chassis along with the initial alignment of the cameras and mirrors in the 3F region. The rework of the video mux and distribution system has also begun this month. A considerable amount of effort has gone into planning and documenting the AMS/Video system. One goal is to have all of the 3F region viewers completely operational with video to the control room by July 4 break.

Machine Protection System planning is now in full swing. The design review will be held the week of July 8. A very simple and sound approach is being taken; in principle each system interfaces to a single opto-isolated VME board which intern generates allowable MACHINE MODES. These boards are then connected to a master which generates the allowable BEAM MODES. This in turn feeds the Drive Laser Pulse Controller (DLPC) which controls the electron source.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• All four core plate sets for the IR machine passed gap inspection, having uniformity variation of no greater than .0005 inch (1/4 human hair) The cores are now at MEI going through inspection. The first unit will go through a pre assembly of al its parts to verify bolt and part fit-up.

• The first GW unit passed pre assembly with a few fastener lengths requiring change.

• MEI performed a successful prototyping of their Purcell Gap Shim glue down procedure within a freshly constructed clean room. The shims were uniformly glued to the base plate within .0005.

• MEI glued the first sets of Purcell Gap shims to two dipoles. The first was uniform through most of the area but for one .003 bump. They stripped the shims off and are trying again, this time with booties and hair nets in the enclosure during gluing. The other core’s glue-down quality is being evaluated.

• We have a nagging cosmetic problem with the surface of the coils after potting. After removal from the potting mold, within an hour, welt-like ridges start to form. They are .005 high, with white, flake-like spots generate out of the originally flat, clear surfaces. The ridges and spots nominally form above the space between conductors, mostly in the flats where the bundle is squeezed to make the coil fit in the coil gap.

• Our epoxy consultant, Jeffrey Hubrig of Innovation Services tracked down a case where similar welts in the epoxy outer surface along the zone between conductors were solved. The manufacturer out gassed the material at higher temperature for longer times. We tried this process and the use of another batch of epoxy. Both didn’t change. At last, using another, similar epoxy without flexibilizer did work on Coil #14.

• Cores for the Second batch of 6 UV Magnets remain in final machining.

Injector Dipoles (DU/DV)

• At Master Machine, Gap check on the first GV Unit showed uniformity to .00014 inches (Sepecification is .00030. The first GV core was shipped out to Wang NMR on the 28th and the remaining 5 units are ready for gap inspection and shipment on the 2nd of July. GU cores will follow.

• At Wang NMR, all GV and GU coils are potted and ready for insertion in cores.

Arc 180 Degree Dipoles (GY)

• I inspected the first core and all parts that had arrived at Wang NMR on my visit on 6/6/02. The parts look good and not damaged.

• The assembled 2nd Core looks good at Bosma and passed its gap check. I quote from my acceptance memo: "If I look at it the way the wide beam of electrons sees this gap and its field, that is at any one radial set of measurements, the uniformity among measurements is within .0002 inches.

• At Wang NMR, the second GY Coil (of four) is in winding its last layer – much better than first.

• GY Coil Potting fixture now designed. They expect ten-day turn around on its manufacture. Coils will be ready by end of August. We have to work shim glue down in parallel

• The assembled 2nd Core passed its gap check at Bosma. I quote from my acceptance memo: "If I look at it the way the wide beam of electrons sees this gap and its field, that is at any one radial set of measurements, the uniformity among measurements is within .0002 inches. It was received by Wang from Bosma.

• At Wang NMR, the second GY Coil (of four) was completed. They improved and reworked the tool substantially – they have instant feedback to a skilled tooling machinist. The third GY Coil (of four) is half wound.

• GY Coil Potting fixture is in fabrication.

• GG Coils were potted.

• Engineering Change Orders that correct some problems with the water fittings of the GG Coil leads were transmitted to Wang.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• Process Equipment Co. (PECo) is nearing completion of the GQ Cores, aiming for a July 9 inspection date with completion of the GX cores lagging by two weeks.

• We approved a change in the PECo Contract to Revision A of the drawings for both GQ and GX cores that add the shim system to the pole tip edges that makes them usable throughout the required energy range.

• At Wang NMR, seven of the eight coils for the GQ Dipoles ware now wound.

Multipole Magnets

3 inch quad (QX)

• The last shipment of QX magnets (for the UV branch) was received and measured.

• We have finally resolved our magnet measurement issues and production measurement of all the QX quads resumed on an 18 hr per day basis, finally completing all QX quads for both the IR and the UV machine.

• We assigned all the QX quadrupoles to their respective locations in the IR machine and boxed up the remaining 23 quads for the UV branch for assignment later.

Trim Quad (QT)

• Milhous Control of Virginia finally got two production units to us.

• At last a potential solution to the four required vertical correctors in the arcs. (GCs) Tom Hiatt's analysis shows that the Trim Quad (QT) is a perfect base for an accurate vertical dipole corrector if the main top and bottom coils are run at 1.6 A and we fill in the vacant corners of the Panofsky coils of the quad with separate coils run at about the same current density and in the proper vertical corrector dipole sense.

• The only downside is that to power the upper/lower coils in this configuration, where one has subtracted current (from the quadrupole current) and the other has current addition, we need two power supplies per corrector magnet that have to be ganged to run in parallel at the same current,

• We are now analyzing the use of the integral GC vertical corrector at several levels of quadrupole operation in anticipation of the change request required for this innovation

Sextupole (SF)

• DULY Research supplied us with check prints of the entire magnet. We also made a last minute switch in the pinning/bolting configuration that went to single 5/8 bolts and 3/8 pins that have larger distances between them so that the structure is not compromised.

Octupole (OT)

• Work started up on this magnet. We will have DULY design this magnet at a simpler level because we will probably make or assemble much of it in-house.

Beam Line and Vacuum

• The bids for the ARC Chambers were received. The technical evaluation team met and found a clear acceptable bid at a good price and 10 week delivery. The procurement was placed with Master Machine of Newport News, VA

• The X Chamber is coming together well in the Shop.

• The drawings for the recirculation region in front of the Cryomodule were not signed off because of some problems with the dump and some reversed antenna leads on a BPM. However the positions of stands and cartridges were checked and the girders can now be released for construction because they are verified to be identical to the 2F regions 8 girders.

• Design of the four chambers for the optical chicane region are nearly complete and went through review of the Tera Hz port.

• Design of the regions around the wiggler continues

WBS 11 (Optics):

We have completed fabrication of the second optical cavity assembly and test stand, and are awaiting receipt of the first parts (leadscrew fixtures, etc), due later this month. The vendor for the OCMMS beamline parts has informed us they will deliver the parts nearly one month late. While there is sufficient schedule float to not be an issue, the same vendor is making a sizeable portion of the optical cavity internals, and we do want to know whether they will make their schedule. The OCMMS external mirrors were inspected and found to meet specs. The drift distance from the external OCCMS mirror to the center of the wiggler was determined, and will be considered in making the optical layout. We received a budgetary estimate for optical tables to support the OCMMS hardware. Given the cost, we are evaluating the use of simpler breadboard-style optical benches. The first tests using Ga:In eutectic as a heat conduction compound showed that a gap of 0.050" was too large. We are going to try again with a much narrower gap. So, far, tests of the mirror mounted with the improved In:Ag brazing technique continue to look very good.

We anticipate brazing of the first article 50kW optical beam dump within the next two weeks. This is a continued slip from our original date, but this item has considerable schedule float. Given the unwillingness of on vendor of metal foam to sell us their material, we are checking with other vendors. As an alternative to the black copper oxide coating we are considering, the anticipated backscattered radiation level using Molectron's flame-sprayed coating (which we know worked well for our 15 kW beam dumps) was modeled. The most probable number of reflections is ~ 4, and with the manufacturer's value for the reflectance (4%) we make our backscatter specification.

Our compact spectrometer arrived this month. It will be used as a pulsewidth monitor (for our ultrafast laser and the compressed output of the drive laser) and for spectroscopic analyses in the near IR. The ultrafast laser oscillator shipped on May 31st and we received it early in June. We received parts for the prototype backplane cooled mirror mount and conveyor chain that will be used for routing of the water lines in the mirror cassettes in the User Labs. Some brazing work remains on the backplane cooled mount, and then it will be checked in the Mirror Test Stand. We are also performing some temporary modifications to the UL 3 mirror cassette in order to perform particulate studies. These are needed to certify our UHV lubrication procedures.

We've signed the drawings for the optical cavity vacuum vessels, and the procurement process is underway. We anticipate receipt by mid-August. The Lab machine shop says we are still on schedule (6/24) on delivery of a portion of the cavity internals. We received the 4" UHV gate valves that isolate the vacuum vessels from the rest of the beamline.

The ultrafast laser oscillator was installed this month. It is set up for 1 ps operation; the service engineer will return and set it up for ultrashort pulse operation. We have begun defining the research plan for damage testing UV optics with this laser.

We received the main mounting plates and various fabricated hardware for the optical cavities. It is in our Machine Shop, going through QC. We anticipate receiving the final QC on the OCMMS beamline parts next week. The OCMMS support stands are designed, and we are getting quotes for their fabrication and the optical tables that rest on them. The first prototype, compact 50kW beam dump is being brazed. The optical cavity vacuum vessels are being procured.

We've received the prototype backplane cooled mirror holder and a test mirror. This will be checked on the mirror test stand in the near future. The designs for the mirror cassette modifications and the turning cassettes are underway.

The ultrafast laser oscillator was converted to femtosecond operation, but not without some problems. Since they were associated with the installation process, we think they are one time occurrences, and we should get dependable operation from this laser. The pulsewidth is less than 80 fs, and we are in the process of checking its other performance parameters. Some additional optics were ordered for the UV optics damage test projects.

Aerospace Corporation June 2002 Progress Report

The laser micromachining station definition phase is coming to a close. Specific system requirements are being collected from the various vendors (such as motion control). A layout of the optical train that feeds the micromachining station has been designed and will be modeled for sensitivity testing using ZmaxTM software. ZmaxTM optical diagnostics software was procured. Aerotech Corporation software engineers visited Aerospace to discuss their view of the CAD to CAM software feature that is required. A three axis MasterCAMTM software has been procured for the CAM/toolpath generation. This procurement follows a 45 day evaluation period where the software was tested for toolpath generation from 3D models drawn with the CAD portion of MasterCAM, SolidWorksTM and AutoCADTM . The proposed vendor (Aerotech) for the motion control system was given some toolpath files for testing in May. This was reported in last month’s progress report. The results of these tests show that additional evaluation into the hardware will be needed before proceeding with procuring the motion control hardware. The current design calls for XY motion of at least 12 inches and a Z motion of at least 3 inches. The Z axis motion limit will essentially limit the types of part shapes that can be processed. The XY lengths are dictated by the surface flatness requirements imposed by high accuracy XY stage vendors like Aerotech (to meet their resolution and speed specification). These surface flatness requirements (12 microns over 1 meter) cannot be met by vibration isolation table vendors like Newport Corporation. Procuring of an all-granite polished table was discussed and there is a significant issue with weight and transportability. The design team is exploring a hybrid (granite/honeycomb structure) surface that could meet the motion control vendor needs. Discussions were also held on finding a non mechanical approach to laser power control. Discussions with various acousto-optic vendors and electro-optic vendors were carried out and an initial design has been proposed.